OpenStudy (anonymous):

can someone please explain the steps for:

OpenStudy (anonymous):

\[\frac{ 14 }{ 18 }=\frac{ x }{ 63 } = x=49\]

OpenStudy (akashdeepdeb):

What is the question? :)

OpenStudy (anonymous):

What is the solution for x in the proportion? Explain your steps.

OpenStudy (akashdeepdeb):

So you are given 14/18=x/63 ?

OpenStudy (anonymous):

yes.

OpenStudy (akashdeepdeb):

Okay so all you need to do in these types of questions to find x is \[\frac{ 14 }{ 18 } = \frac{ x }{ 63 }\] Take 63 to the other side \[\frac{ 14 * 63 }{ 18 }\] = x Now divide numerator by a common factor such as 9!! :D \[\frac{ 14 * 7 }{ 2 }\] = x Now solve completely. \[\frac{ 7*7 }{ 1 }\] = x So x=49 !! :D Understood? :)

OpenStudy (anonymous):

Also just basically cross multiply, which is what they did

OpenStudy (anonymous):

thank you soooo much! I understand :)

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify \[\sqrt{320}\] and provide the simplified form.

OpenStudy (akashdeepdeb):

Post a new question! :)